# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ 4.2 Convergence Analysis
By (18) and the strict convexity of the function $F(x)$ (cf. Assumption 6), Proposition 4.9 is proved.
Proposition 4.11 (Finite-time Termination).
Take any initial restriction parameters $\epsilon_{i}^{0}>0$, any initial finite or empty sets $Y^{0}_{i}\subset Y_{i}$, and any reduction parameter $r>1$. Under Assumptions 1, 2, 4, 5, 6, the distributed cutting-surface consensus algorithm terminates finitely and obtain solutions for all agents that satisfy consensus and zeroth-order stopping conditions to certain tolerances.
Proof 4.12.
By Proposition 4.9, we obtain the convergence result of the distributed cutting-surface consensus algorithm. To prove the finite-time termination of the algorithm, we only need to exclude three possible infinite cases (Case I-III), which are presented in Figure 1.
1) Inexistence of infinite Case I: By Assumption 1, there exists at least one interior point $\hat{x}$ for the (${\rm DRCP}$), we have
where $\epsilon_{i}^{s}$ is a positive scalar parameter for agent $i$ satisfying $0<\epsilon_{i}^{s}\leq\min_{y_{i}\in Y_{i}}{-g_{i}(\hat{x},y_{i})}$. Assume that the (${\rm ADRCP^{k}}$) is unsolvable at the $k$-th iteration, i.e., $\mathcal{F}^{k}=\left\{x\in X|g_{i}(x,y_{i})\leq-\epsilon^{k}_{i},\ \forall y_{i}\in Y_{i}^{k},\ i=1,...,m\right\}=\emptyset$. Then, after at most $K_{1}=\max_{i\in\mathcal{V}}\lceil\log_{r}\epsilon^{k}_{i}/\epsilon^{s}_{i}\rceil$ updates of restriction parameters $\epsilon^{k}_{i}$, the relations $0<\epsilon_{i}^{K_{1}}\leq\epsilon_{i}^{s}$ and $Y_{i}^{K_{1}}\subset Y_{i}$ hold. According to this result and (19), we have $\mathcal{F}^{K_{1}}\supset\hat{\mathcal{F}}\neq\emptyset$. Therefore, after at most $K_{1}$ executions of Case I in Algorithm 1, the (${\rm ADRCP^{k}}$) becomes solvable.
2) Inexistence of infinite Case II: Referring to the proof idea in [32], we prove that the infinite Case II does not exist for any agent in the loop of Algorithm 1. For any $i\in\mathcal{V}$, we first consider a sequence of solutions of this agent in the (${\rm ADRCP^{k}}$). Since $X$ is a compact set and Proposition 4.9 holds, we can select a converging subsequence for agent $i$ satisfying $x^{k}_{i}\rightarrow\widetilde{x}$ as $k\rightarrow\infty$. Assume that at the $k$-th iteration, the point $x^{k}_{i}$ does not satisfy its local constraints $\mathcal{F}_{i}$ and $\overline{y}^{\max}_{i}$ is obtained through solving the (${\rm LLP_{i}}$). Thus, the following relation holds:
$g_{i}(x^{k}_{i},\overline{y}^{\max}_{i})>0.$
In step 15 of Algorithm 1, the point $\overline{y}^{\max}_{i}$ is added into the finite set of agent $i$. Hence, for any $l>k\geq 0$, it follows that $\overline{y}^{\max}_{i}\in Y^{l}_{i}$. In Section 3.1, the DPG algorithm guarantees the local feasibility of the solution due to the adoption of the Euclidean projection. Therefore, for any $l,k\ (l>k\geq 0)$, it holds that
$g_{i}(x^{l}_{i},\overline{y}^{\max}_{i})\leq-\epsilon^{l}_{i}.$
$\exists\delta>0\ \ g_{i}(x,\overline{y}^{\max}_{i})\leq-\epsilon^{l}_{i}/2<0\ \ \|x-x^{l}_{i}\|<\delta\ \ \forall l>k\geq 0$ (20)
As a result of $x^{k}_{i}\rightarrow\widetilde{x}$, for any $\delta>0$ there is
$\exists K_{2}:\ \|x^{l}_{i}-x^{k}_{i}\|<\delta\ \ \forall l,k:\ l>k\geq K_{2}$ (21)
Combine with (20) and (21), we can conclude that,
$\exists K_{2}:\ g_{i}(x^{k}_{i},\overline{y}^{\max}_{i})\leq-\epsilon_{i}^{l}/2<0\ \ \forall k:\ k\geq K_{2}$
Therefore, we can conclude that for any agent $i\in\mathcal{V}$, after at most $K_{2}$ iterations of Algorithm 1, the approximate optimal point $x^{K_{2}}_{i}$ of the agent $i$ obtained by solving the (${\rm ADRCP^{K_{2}}}$) satisfies its local constraints $\mathcal{F}_{i}$ in the $({\rm DRCP})$.
3) Inexistence of infinite Case III: Suppose that the (${\rm ADRCP^{k}}$) is solvable at the $k$-th iteration of Algorithm 1. By using DPG algorithm equipped with distributed finite-time termination algorithm, each agent can obtain an approximate optimal solution $x^{k}_{i}$ of the (${\rm ADRCP^{k}}$) (cf. Corollary 3.4). Moreover, for any agent $i\in\mathcal{V}$, we have
$\|x^{k}_{i}-x^{k,*}\|=\delta^{k}_{i}$ (22)
where $k\geq 0$. $\delta^{k}_{i}$ is a positive scalar satisfying $0\leq\delta^{k}_{i}\ll\min\left\{\epsilon_{4},\epsilon_{5},\epsilon_{6}\right\}$ as the DPG algorithm is fully iterated by setting small parameters $\epsilon_{1},\epsilon_{2},\epsilon_{3}$, where $\max\left\{\epsilon_{1},\epsilon_{2},\epsilon_{3}\right\}\ll\min\left\{\epsilon_{4},\epsilon_{5},\epsilon_{6}\right\}$. Then, we can set $x^{k}_{i}=x^{k,*}+\delta^{k}_{i}\cdot v^{k}_{i}$, where $v^{k}_{i}$ is a unit vector equals to $({x^{k}_{i}-x^{k,*}})/\|x^{k}_{i}-x^{k,*}\|$. For any two agents $i,j\in\mathcal{V}$, suppose that they update $z_{i}^{k_{i}},z_{j}^{k_{j}}$ by Step 6 at the $k_{i}$-th and $k_{j}$-th iterations of Algorithm 1, respectively. Therefore, we have
$\|z_{i}^{k_{i}}-z_{j}^{k_{j}}\|$ $=\|x_{i}^{k_{i}}-x_{j}^{k_{j}}\|$ (23)
$=\|x^{k_{i},*}+\delta^{k_{i}}_{i}\cdot v^{k_{i}}_{i}-x^{k_{j},*}-\delta^{k_{j}}_{j}\cdot v^{k_{j}}_{j}\|$
$\leq\|x^{k_{i},*}-x^{k_{j},*}\|+\|\delta^{k_{i}}_{i}\cdot v^{k_{i}}_{i}-\delta^{k_{j}}_{j}\cdot v^{k_{j}}_{j}\|$
$\leq\|x^{k_{i},*}-x^{k_{j},*}\|+\delta^{k_{i}}_{i}+\delta^{k_{j}}_{j}.$
where the two inequalities hold based on the triangle inequality. Next, we discuss the limit on both sides of (23) as $k_{i}\rightarrow\infty$ and $k_{j}\rightarrow\infty$. According to the convergence of $\left\{x^{k,*}\right\}_{k}$ (cf. Proposition 4.9), it follows that
$\mathop{\lim}\limits_{k\rightarrow\infty}\|x^{k,*}-x^{*}\|=0.$
By the triangle inequality, there is for any $k_{i},k_{j}\geq 0$,
$\|x^{k_{i},*}-x^{k_{j},*}\|\leq\|x^{k_{i},*}-x^{*}\|+\|x^{k_{j},*}-x^{*}\|.$ (24)
Take the limit on both sides of the inequality (24) as $k_{i}\rightarrow\infty$ and $k_{j}\rightarrow\infty$, it holds that
$\mathop{\lim}\limits_{k_{i},k_{j}\rightarrow\infty}\|x^{k_{i},*}-x^{k_{j},*}\|$ (25)
$\quad\leq\mathop{\lim}\limits_{k_{i}\rightarrow\infty}\|x^{k_{i},*}-x^{*}\|+\mathop{\lim}\limits_{k_{j}\rightarrow\infty}\|x^{k_{j},*}-x^{*}\|=0.$
Next, take the limit on both sides of the inequality (23) and combine the relation (25), we can get:
$\mathop{\lim}\limits_{k_{i},k_{j}\rightarrow\infty}\|z_{i}^{k_{i}}-z_{j}^{k_{j}}\|=\mathop{\lim}\limits_{k_{i},k_{j}\rightarrow\infty}\|x^{k_{i}}_{i}-x^{k_{j}}_{j}\|\leq\mathop{\lim}\limits_{k_{i},k_{j}\rightarrow\infty}\delta^{k_{i}}_{i}+\delta^{k_{j}}_{j}$
@@
Since $\delta^{k_{i}}_{i}+\delta^{k_{j}}_{j}\ll\epsilon_{4}$ holds for any $k\geq 0$, the consensus stopping condition (9a) can be met in finite iterations of Algorithm 1. Similarly, we can verify that the last two conditions in the stopping criterion (9) can also be satisfied within finite iterations. Therefore, we can obtain that Algorithm 1 terminates finitely and the solutions for all agents satisfy consensus and zeroth-order stopping conditions to certain tolerances.
@@
[…]
Proof 4.14.
[…]
@@
Under Assumptions 1, 2, 4, 5, 6, the distributed cutting-surface consensus algorithm terminates finitely and generates locally feasible solutions for all agents satisfying consensus and zeroth-order stopping conditions to certain tolerances. This holds for any initial restriction parameters $\epsilon_{i}^{0}>0$, any initial finite or empty sets $Y^{0}_{i}\subset Y_{i}$, and any reduction parameter $r>1$.
Proof 4.16.
This result follows from Proposition 4.11 and 4.13.
@@ 5 Numerical case studies
In this section, a numerical case is presented to validate the effectiveness of the cutting-surface consensus algorithm (i.e., Algorithm 1). Firstly, we analyze and verify the finite-time termination of Algorithm 1 and confirm the effectiveness of the distributed termination algorithm developed in this paper by comparing it with the finite-time consensus termination algorithm in [37]. Then, we verify the local feasibility of the solutions obtained by Algorithm 1 and compare it with the state-of-the-art methods [29, 30]. Finally, we analyze the impacts of the restriction parameter $\epsilon^{0}_{i}$ and the reduction parameter $r$ on Algorithm 1. The implementation is carried out in MATLAB Version 9.5.0.944444 (R2018b, win64) and runs on an Intel(R) Core (TM) i7-7700HQ CPU @ 2.80GHz, 256GB terminal server.
Figure 3: Network Topology: time-varying unbalanced directed graphs, where the black and blue arrows denote the communication among agents at time slot $t$ and $t+1$. $\mathcal{G}(t:t+1)$ is strongly connected.
@@
[…]
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\forall y_{i}\in[-1,1]\quad i\in\mathcal{V},$
@@
Table 2: Parameters of The Problem
@@ 5.1 Analysis of Finite-Time Termination
Figure 4: Iterative process of Algorithm 1. The blue solid line indicates the evolution of $F=\sum^{m}_{i=1}f_{i}(z^{k}_{i})$ with respect to iteration $k$. The red dashed line presents the global optimal value of the problem (26), where $x^{*}=[0,\sqrt{7}/4]^{\top}$. The horizontal coordinates of the black and red dotted-dashed lines represent the number of iterations required by the finite-time consensus termination algorithm in [37] and the termination algorithm developed in this paper, respectively.
@@
For Algorithm 1, we initiate the restriction parameter $\epsilon^{0}_{i}=100$ and the initial set $Y^{0}_{i}=\emptyset$ for each $i\in\mathcal{V}$, along with the reduction parameter $r=10$. For Algorithm 2, we set the termination parameters as $\epsilon_{4}=0.1$, $\epsilon_{5}=0.1$ and $\epsilon_{6}=0.1$. Moreover, for the DPG algorithm equipped with the distributed termination method, we set the stepsize as $\alpha(t)=\frac{1}{\sqrt{t}}$, and the termination parameters as $\epsilon_{1}=10^{-2}$, $\epsilon_{2}=10^{-6}$, $\epsilon_{3}=10^{-6}$.
@@
For this example, the distributed cutting-surface consensus algorithm terminates after eight iterations. Figure 4 shows the iterative process of Algorithm 1, from which it can be seen that with the iteration of Algorithm 1, the global objective $F\!=\!\sum^{m}_{i=1}f_{i}(z^{k}_{i})$ converges to the optimal value $F^{*}$ of the problem (26). Note that in the first three iterations of Algorithm 1, since there exists at least an agent $i\in\mathcal{V}$ that fails to execute Step 6 in Algorithm 1, its candidate point $z^{k}_{i}$ satisfies $\|z^{k}_{i}\|\!=\!\|z^{0}_{i}\|\!=\!+\infty$. Therefore, we have $\sum^{m}_{i=1}f_{i}(z^{k}_{i})=+\infty$.
@@
In addition, we discuss the iterative process of the DPG algorithm when the local constraints $\mathcal{F}^{k}_{i}$ of all agents in the (${\rm ADRCP^{k}}$) are nonempty. Figure 5 presents that the DPG algorithm reaches convergence when the (${\rm ADRCP^{k}}$) is solvable, i.e. $\mathcal{F}^{k}=\cap^{m}_{i=1}\mathcal{F}^{k}_{i}\neq\emptyset$. Figure 6 illustrates that the DPG algorithm fails to reach consensus and optimality when the (${\rm ADRCP^{k}}$) is unsolvable, i.e. $\mathcal{F}^{k}=\emptyset$. As described in Section 4.1, to ensure finite-time termination of Algorithm 1, we set the maximum number of time slots $T=10^{7}$ to force the termination of the DPG algorithm for this situation.
@@
Figure 5: Iterative process of the DPG algorithm, when the (${\rm ADRCP^{k}}$) is solvable ($\epsilon^{k}_{i}=0.1$ and $Y^{k}_{i}=[1]$ for all agents). The six solid lines indicates the evolution of $\|x^{k}_{i}(t)-x^{k,*}\|$ with the time slot $t$, where $x^{k}_{i}(t)$ is agent $i$’s state in the time slot $t$ of the DPG algorithm. The red dashed line presents the value $\|x^{k}_{i}(t)-x^{k,*}\|=0$. The horizontal coordinates of the black and red dotted-dashed lines represent the number of iterations required by the finite-time consensus termination algorithm in [37] and the proposed termination algorithm, respectively. The zoomed-in part shows our optimality gap is smaller than that achieved by [37].
@@
Figure 6: Iterative process of the DPG algorithm, when the (${\rm ADRCP^{k}}$) is unsolvable ($\epsilon^{k}_{i}=5$ and $Y^{k}_{i}=[1]$ for all agents). The six solid lines indicates the evolution of $\|x^{k}_{i}(t)-\overline{x}^{k}(T)\|$ with the time slot $t$, where $x^{k}_{i}(t)$ denotes agent $i$’s state in the time slot $t$ of the DPG algorithm, $T=10^{-7}$ and $\overline{x}^{k}(T)=\frac{1}{m}\sum_{i=1}^{m}x^{k}_{i}(T)$.
@@
Moreover, we compare our termination algorithm with the finite-time consensus algorithm in [37]. In solving the (DRCP), the finite-time consensus algorithm achieves termination in the fourth iteration of Algorithm 1, as shown in Figure 4. However, the resulting solution is not the approximate optimal solution for the (DRCP). In contrast, our distributed termination algorithm not only ensures the consensus of the solutions but also guarantees their approximate optimality upon termination of Algorithm 1. Besides, when solving the (${\rm ADRCP^{k}}$), our solution is more accurate than that of the finite-time consensus algorithm [37] under the same consensus accuracy $\epsilon_{1}$. This stems from the introduction of the zeroth-order stopping conditions (4b) and (4c), enabling the DPG algorithm to generate a more approximate optimal solution for each agent, as shown in Figure 5.

### v3

@@ 3.4 Convergence Analysis
By (23) and the strict convexity of the function $F(x)$ (cf. Assumption 6), it can be obtained that $x^{k,*}\rightarrow x^{*}$.
Combining the convergence result of the sequence $\{x^{k,*}\}_{k}$ with (19), it can be obtained that for a given $\epsilon_{4}>0$,
$\exists\overline{k}_{i},\overline{k}_{j}:\|z_{i}^{\overline{k}_{i}+1}-z_{j}^{\overline{k}_{j}+1}\|\leq\epsilon_{4},\ \forall i,j\in\mathcal{V}.$
@@
Hence, the consensus stopping condition (10a) can be met in finite iterations of Algorithm 2. Similarly, we can verify that the last two conditions in the stopping criterion (10) can also be satisfied within finite iterations. Therefore, we can obtain that Algorithm 2 terminates finitely and the solutions for all agents satisfy consensus and zeroth-order stopping conditions to certain tolerances.
@@
Suppose Assumptions 1, 2, 46 are satisfied. The distributed cutting-surface consensus algorithm generates local feasible solutions for the ($\rm DRCP$}), i.e., $z_{i}^{k+1}\in\mathcal{F}_{i},\ \forall i\in\mathcal{V}$.
Proof 3.12.
In Algorithm 2, since $({\rm LLP}_{i})$ is solved at optimality, it can be obtained that for any $i\in\mathcal{V}$, each term of the sequence $\{z_{i}^{k+1}\}_{k}$ either belongs to the set $\mathcal{F}_{i}$ or satisfies $\|z_{i}^{k+1}\|=+{\rm Inf}$. By Proposition 3.9, we have that Algorithm 2 terminates finitely, and $\|z_{i}^{k+1}\|<+{\rm Inf}$ for all $i\in\mathcal{V}$. Therefore, when Algorithm 2 terminates, the obtained point $z_{i}^{k+1}$ satisfies $z_{i}^{k+1}\in\mathcal{F}_{i},\ \forall i\in\mathcal{V}$.
Remark 3.13.
In solving the ($\rm{ADRCP}^{\it k}$), Algorithm 1 cannot provide an exact optimal solution that simultaneously satisfies the constraints of all the agents in finite time, but can only find a solution that satisfies its own constraints of each agent in finite time. Thus, the local feasibility of the solutions to the ($\rm{ADRCP}^{\it k}$) leads directly to the fact that the proposed distributed cutting-surface consensus algorithm does not return in finite time solutions that are feasible for the simultaneous satisfaction of semi-infinite constraints for all agents.
@@
For each agent $i\in\mathcal{V}$, take any initial restriction parameter $\epsilon_{i}^{0}>0$, any initial finite or empty set $Y^{0}_{i}\subset Y_{i}$, and any reduction parameter $r>1$. Under Assumptions 1, 2, 46, the distributed cutting-surface consensus algorithm terminates in a finite number of iterations and generates locally feasible solutions for all agents satisfying consensus and zeroth-order stopping conditions to certain tolerances.
Proof 3.15.
This result follows from Propositions 3.9 and 3.11.
@@ 4 Numerical case studies
In this section, a numerical case is presented to validate the effectiveness of the cutting-surface consensus algorithm (i.e., Algorithm 2). Firstly, we analyze and verify the finite-time termination of Algorithm 2 and confirm the effectiveness of the distributed termination algorithm developed in this paper by comparing it with the termination algorithm in [38]. Then, we verify the local feasibility of the solutions obtained by Algorithm 2 and compare it with the state-of-the-art methods in [26, 27]. Finally, we analyze the impacts of the restriction parameter $\epsilon^{0}_{i}$ and the reduction parameter $r$ on Algorithm 2. The implementation is carried out in MATLAB Version 9.5.0.944444 (R2018b, win64) and runs on an Intel(R) Core (TM) i7-7700HQ CPU @ 2.80GHz, 256GB terminal server.
Figure 2: Network Topology: time-varying unbalanced directed graphs, where the black and blue arrows denote the communication among agents at time $t$ and $t+1$. $\mathcal{G}(t:t+1)$ is strongly connected.
@@
$\ \text{s.t.}\ g_{i}(x,y_{i})=(x(1)-p_{i})^{2}+2y_{i}x(2)-{y_{i}}^{2}\!-\!1\!\leq\!0$
$\qquad\qquad\qquad\qquad\qquad\qquad\ \ \forall y_{i}\in[-1,1]\quad i\in\mathcal{V},$
@@
Table 2: Parameters of The Numerical Problem
@@ 4.1 Analysis of Finite-Time Termination
Figure 3: Iterative process of Algorithm 2. The blue solid line indicates the evolution of $F=\sum^{m}_{i=1}f_{i}(z^{k}_{i})$ with respect to iteration $k$. The red dashed line presents the global optimal value of the problem (24), where $x^{*}=[0,\sqrt{7}/4]^{\top}$, $F^{*}\approx 38.687746$. The horizontal coordinates of the black and red dotted-dashed lines represent the number of iterations required by the finite-time consensus termination algorithm in [38] and the termination algorithm developed in this paper, respectively.
@@
For Algorithm 2, we initialize the restriction parameter $\epsilon^{0}_{i}\!=\!100$ and the initial set $Y^{0}_{i}\!=\!\emptyset$ for each $i\in\mathcal{V}$, along with the reduction parameter r=10r\!=\!10. For Algorithm 1, we set the stepsize as $\alpha(t)=\frac{1}{\sqrt{t}}$, and the termination parameters as $\epsilon_{1}\!=\!10^{-2}$, $\epsilon_{2}\!=\!10^{-6}$, $\epsilon_{3}\!=\!10^{-6}$. For Algorithm 3, we set the termination parameters as $\epsilon_{4}\!=\!0.1$, $\epsilon_{5}\!=\!0.1$ and $\epsilon_{6}\!=\!0.1$. In addition, in the execution of Algorithm 1, the Euclidean projection point for each agent is obtained by solving the sub-optimization problem introduced in Remark 4, using the fmincon function in MATLAB.
@@
For this example, the distributed cutting-surface consensus algorithm terminates after eight iterations. Figure 3 illustrates the global objective value $F\!=\!\sum^{m}_{i=1}f_{i}(z^{k}_{i})$ at different iterations of Algorithm 2, from which it can be seen that the global objective $F$ converges to the optimal value $F^{*}$ of the problem (24) as the iterations progress. Note that in the first three iterations of Algorithm 2, since there exists at least an agent $i\in\mathcal{V}$ that fails to execute Step 13 in Algorithm 2, its candidate point $z^{k}_{i}$ satisfies $\|z^{k}_{i}\|\!=\!\|z^{0}_{i}\|\!=\!+\infty$. Therefore, we have $\sum^{m}_{i=1}f_{i}(z^{k}_{i})=+\infty$.
@@
In addition, we discuss the iterative process of the DPG algorithm when the local constraints $\mathcal{F}^{k}_{i}$ of all agents in the ($\rm ADRCP^{k}$}) are nonempty. Figure 4 presents that the DPG algorithm reaches convergence when the ($\rm ADRCP^{k}$}) is feasible, i.e. $\mathcal{F}^{k}=\bigcap^{m}_{i=1}\mathcal{F}^{k}_{i}\neq\emptyset$. Figure 5 illustrates that the DPG algorithm fails to reach consensus when the ($\rm ADRCP^{k}$}) is infeasible, i.e. $\mathcal{F}^{k}=\emptyset$. As described in Section 3.3, to ensure finite-time termination of Algorithm 2, we set the maximum iteration time $T=10^{7}$ to force the termination of the DPG algorithm for this situation.
@@
Figure 4: Iterative process of the DPG algorithm, when the ($\rm ADRCP^{k}$}) is feasible ($\epsilon^{k}_{i}=0.1$ and $Y^{k}_{i}=[1]$ for all agents). The six solid lines indicates the evolution of $\|x^{k}_{i}(t)-x^{k,*}\|$ with respect to $t$, where $x^{k}_{i}(t)$ is agent $i$’s estimate generated by DPG algorithm at time $t$. The red dashed line presents the value $\|x^{k}_{i}(t)-x^{k,*}\|=0$. The horizontal coordinates of the black and red dotted-dashed lines represent the number of iterations required by the finite-time consensus termination algorithm in [38] and the proposed termination algorithm, respectively. The zoomed-in part shows our optimality gap is smaller than that achieved by [38].
@@
Figure 5: Iterative process of the DPG algorithm, when the ($\rm ADRCP^{k}$}) is infeasible ($\epsilon^{k}_{i}=5$ and $Y^{k}_{i}=[1]$ for all agents). The six solid lines indicates the evolution of $F(x^{k}_{i}(t))$ with respect to $t$, where $x^{k}_{i}(t)$ denotes the agent $i$’s estimate generated by DPG algorithm at time $t$.
@@
Moreover, we compare our termination algorithm with the finite-time consensus algorithm in [38]. In solving the (DRCP), the finite-time consensus algorithm achieves termination in the fourth iteration of Algorithm 2, as shown in Figure 3. However, the resulting solution is not the approximate optimal solution for the (DRCP). In contrast, our distributed termination algorithm not only ensures the consensus of the solutions but also guarantees their approximate optimality upon termination of Algorithm 2. Moreover, when solving the ($\rm ADRCP^{k}$}), the optimality gap of our solution is smaller than that achieved by the finite-time consensus algorithm in [38] under the same consensus accuracy $\epsilon_{1}$. This stems from the introduction of the zeroth-order stopping conditions (4b) and (4c), enabling the DPG algorithm to generate a more approximate optimal solution for each agent, as shown in Figure 4.